Please calculate the amount of energy it will take to change 25.0 Grams of Water from -6 C to changing it all to vapor at sea level. Show all your work for credit.




Please calculate the amount of energy it will take to change 25.0 Grams of Water from -6 C to changing it all to vapor at sea level. Show all your work for credit.

heat of fusion of water = 334 J/g

heat of vaporization of water = 2257 J/g

specific heat of ice = 2.09 J/g·°C

specific heat of water = 4.18 J/g·°C
1
Expert's answer
2013-02-28T10:35:51-0500
Heating 25 g of ice from -6 to 0 C needs next amount of energy:

E = 25 g*2.09 J/g·°C * 6 = 313.5 J

Heating 25 g of water from 0 C to 100 C needs next amount of energy:

E = 25 * 4.18 J/g·°C * 100 = 1045000 J

Amount of energy for changing from solid to liquid phase is:

E = 25 * 334 = 8350 J

Sum of all energy is 1053663.5 J or 1053.7 kJ